could you please help me solve this question?
The probability that a married couple pays for their honeymoon is 0.56. You randomly select 5 couples and ask each if they paid for their honeymoon themselves. Find the probability that exactly 4 of the couples said that they paid for their honeymoon.

Answer by rothauserc(4718)

This problem requires the use of the binomial probability distribution
:
We are given p = 0.56, q = 1 - 0.56, k = 4, and n = 5, then
:
P (k = 4) = 5C4 * (0.56)^4 * (0.44)^1 =
:
(5! / (4! * (5-4)!) * (0.0983) * (0.44) = 0.2164
:
*****************************************************
The probability that exactly 4 couples paid for their
honeymoon is approx 0.22
*****************************************************
: